Skip to main content

Rena update from Maritime NZ (217)

Friday, 27 January 2012 11:00 a.m.

Salvage

  • Svitzer salvors yesterday removed 8 empty containers from the bow section of the Rena, by sliding the containers into the water. Of these, 6 were recovered from the water by theSmit Borneocrane, and two have sunk to the reef - these have had buoys attached to them and will be recovered once the weather conditions allow.
  • Salvors also removed 39 bags of milk powder yesterday, emptying one complete container.
  • Weather conditions are not favourable for salvage operations today, with winds of up to 30 knots and swells of up to 2m. The weather is forecast to remain unsettled over the weekend. The salvage team will take advantage of good weather windows when they can.
  • It is unlikely any container removal will be undertaken today, so salvors will continue with the emptying of containers on board and removal of bags of cargo.

Container and debris recovery

  • The Braemar recovery team has been working in difficult conditions at Bowentown to cut up a container that's partly submerged. It's a challenging task because of strong tidal currents which mean that divers with special underwater cutting equipment can only work at low tide.
  • Shoreline debris clean up is continuing at Matakana Island and on the East Cape.
  • The recovery team has so far processed 497 containers - this comprises 441 that have been lifted off the wreck by the salvors and 56 that have been collected at sea or from beaches by Braemar vessels.
  • Beachgoers are reminded to stay away from any container debris sighted - and asked to please not touch container debris. Members of the public are asked to please phone in any sightings of containers or debris to 0800 333 771. 

Oil spill response

  • A shoreline clean-up assessment team spent yesterday inspecting Blue Gum Bay and Hunters Creek on Matakana Island and found no oil or debris.
  • A second team will be working along East Cape/Te Kaha today - yesterday they located some oiled debris but no oil.
  • Oil spill response teams worked at Tuapiro yesterday and started early this morning to accommodate today's high tide. The tide times mean that teams working at Leisure Island (Motoriki) and Mauāo, will be starting after midday today and working into the evening.
  • An environmental assessment team will be inspecting Rabbit Island today to assess the levels of residual oil there.
  • The total amount of waste collected by the oil spill response team so far is 1020 tonnes.
  • Massey University is now accommodating 26 birds - 25 little blue penguins and one grey faced petrel. One petrel died in care yesterday.
  • One slightly oiled live penguin was found at Maketū Spit yesterday.
  • Te Maunga wildlife site is being inspected to ensure it is ready to be handed back to the owners. Some equipment is still to be moved out.
  • A team of three working on Mōtītī's southern beaches yesterday found no oiled birds or oil, but did pick up some twine and netting.
Rena 27 January